# The interplay of emission geometry and extragalactic magnetic fields in the spectra of ultra-high-energy cosmic rays

### v1

@@ IV.1 The pure-proton scenario
where $\tau$ is the particle’s propagation time. Panel (a) corresponds to $\theta=0^{\circ}$, while panel (b) corresponds to $\theta=80^{\circ}$, both of which show the results for $B=10\;\text{nG}$ and $L_{B}=0.1\;\text{Mpc}$. The analysis of panel (a) reveals that most events reaching Earth have a very small elongation length, close to 1 Mpc. This indicates that the actual distance traveled by these particles is nearly equal to the source–Earth distance, resulting from minimal diffusion. As the arrival energy of these events decreases, the elongation length naturally increases. However, the counts for distances of $d\sim 10^{3}$ Mpc, which is of the same order of magnitude as the distance light would travel during the age of the Universe, are negligible. We conclude that the suppression observed in the lower energy part of the spectrum is not due to the high diffusion of these particles, which would cause the simulation time to reach the age of the Universe, but rather due to magnetic deflection. Unlike panel (a), panel (b), corresponding to $\theta=80^{\circ}$, indicates that a significantly larger number of protons propagated over distances greater than 1 Gpc (not shown in the plot, as the histogram represents the energy of detected particles). Consequently, their propagation times surpass the age of the Universe, preventing them from being detected on Earth.
Figure 3: Event count distribution as a function of the energy ($E$) and the trajectory elongation ($d$). Figure (a) presents the case where $\theta=0^{\circ}$, $B=10.0\;\text{nG}$, $L_{B}=0.1\;\text{Mpc}$ and $R=100\;\text{Mpc}$ while figure (b) shows the results for $\theta=80^{\circ}$, $B=10.0\;\text{nG}$, $L_{B}=0.1\;\text{Mpc}$ and $R=100\;\text{Mpc}$.
IV.2 The case of nuclei
@@
Figure 4: Modification factor obtained for primary iron nuclei. The columns represent different source–Earth distances ($R$, and the rows vary according to the angle between the jet direction and the line of sight ($\theta$). The line styles indicate different magnetic field strengths ($B$) while the line colors correspond to different coherence lengths ($L_{B}$). Vertical lines mark the critical energy, with the color and line style matching the respective coherence length and magnetic field.
Figure 5: Energy distribution for primary nitrogen nuclei, for different arriving compositions. The figures show the results for the case where the jet direction is $\theta=0^{\circ}$, the magnetic strength considered is $B=0.1$ nG and $L_{B}=0.1$ Mpc. The columns vary in source–Earth distance, $R=10,\;40,\;100\;\text{Mpc}$ from left to right.
@@
Fig. 4 shows the modification factor obtained after simulating $10^{5}$ nitrogen. The panels are organized in the same way as it was for protons, with rows corresponding to angles between the jet direction and the line of sight source–Earth and columns to different source–Earth distances. Due to the complexity introduced by nuclear fragmentation, we show energy distributions at Earth separated by particle type, for selected cases. We first analyze the results corresponding to $\theta=0^{\circ}$ (first row). The peak at the highest energies is significantly broader than in the case of proton simulations. This broadening arises because nitrogen nuclei fragment into multiple lighter elements with varying energies, producing a wider spread in the energy distribution of secondary protons arriving at Earth. Figure 5 shows the energy spectra for different nuclei, corresponding to the purple dashed-dotted lines in the first row of Fig.4. These simulations were performed with an angle between the jet direction and the line of sight of $\theta=0^{\circ}$, magnetic field strength $B=0.1\;\text{nG}$, and coherence length $L_{B}=0.1\;\text{Mpc}$. From this figure, we observe that for a source distance of $R=10$ Mpc (first column), the energy range above $10^{20}$ eV is dominated by helium and protons resulting from nitrogen fragmentation. These secondary particles do not travel far enough to undergo photopion production or further nuclear fragmentation. This feature is absent in the spectra for $R=40$ Mpc (second column) and $R=100$ Mpc (third column), where the increased propagation distance enhances the probability of interactions with background photons. Additionally, the purple dashed-dotted curves in the first row of fig. 4 show higher values of $\xi$ at $E\lesssim 6\times 10^{18}$ eV, due to the combination of nitrogen injected at the source with energies below the fragmentation threshold and an insufficient magnetic field strength and/or coherence length to cause larger deflections.
Figure 7: Energy distribution for primary nitrogen nuclei, for different arriving compositions. The panels show the results for the case where the magnetic strength considered is $B=10.0$ nG, the coherence length $L_{B}=10.0\;\text{Mpc}$ and $R=10$ Mpc. The left panel corresponds to a jet inclination of $\theta=40^{\circ}$ and the right one to $\theta=80^{\circ}$.
[…]
@@
With respect to the cases of $\theta=40^{\circ}$ (second row), $\theta=80^{\circ}$ (third row) and the same magnetic field $B=0.1\;\text{nG}$ (dashed-dotted line) discussed above, the curves are dominated by nitrogen emitted with energies below the fragmentation energy threshold. It may be reasonable to assume that the secondary particles produced from nitrogen photodisintegration do not experience enough deflection to reach Earth from misaligned jets with $\theta=40^{\circ}$ or $\theta=80^{\circ}$ in the presence of this magnetic field strength. One can also observe that as the source–Earth distance increases, more energetic particles have higher probability of arriving at Earth for the same parameter space due to the large angular deflection. As a consequence, the maximum energy observed in the spectra is shifted to the right, as expected and discussed in the pure-proton scenario. For a similar reason, the increase of $R$ must reveal an increasing contribution of lighter lower-energy nuclei, although it is expected that the primary nuclei remain dominant in the spectra. Similar conclusions are obtained considering $\theta=80^{\circ}$.
Figure 9: Energy distribution separated by the particle composition for the simulation of iron nuclei as primaries. The figures show the results for the case where the jet direction is $\theta=0^{\circ}$, the magnetic strength considered is $B=0.1$ nG and $L_{B}=1.0$ Mpc. The columns vary in source–Earth distance ($R$), $R=10,\;40,\;100\;\text{Mpc}$ from left to right.
For $B=1\;\text{nG}$ (continuous line), the peak at the highest energies, observed at $\theta=0^{\circ}$ (first row), is narrower for $L_{B}=1\;\text{Mpc}$ and $L_{B}=10\;\text{Mpc}$ compared to the peak for $L_{B}=0.1\;\text{Mpc}$ as a consequence of the loss of lower-energy events, resulting from angular deflections greater than $\Psi/2$. For $\theta=40^{\circ}$ and $\theta=80^{\circ}$, when we compare the results with those corresponding to $B=0.1\text{nG}$ (dashed-dotted lines), we observe that the maximum energy achieved on Earth also increases. This result is in agreement with the conclusions addressed earlier in the pure-proton scenario.
@@
Fig. 8 shows the modification factor resulting after simulating $10^{5}$ iron nuclei as primary particles. We first analyze the results corresponding to $\theta=0^{\circ}$ and $B=0.1\;\text{nG}$. The energy spectra for each particle type for these parameters, assuming $L_{B}=1.0\;\text{Mpc}$, are displayed in fig. 9. For $R=10\;\text{Mpc}$, the peak at $E\sim 10^{19}\;\text{eV}$ observed in fig. 8 is dominated by protons as can be seen in fig 9. In contrast, the peaks at $E\sim 10^{19.5}\;\text{eV}$ and $E\gtrsim 10^{20}\;\text{eV}$ are respectively dominated by helium and iron primaries. For $R=40\;\text{Mpc}$ and $R=100\;\text{Mpc}$, the helium peak approaches the proton peak, as a result of the higher probability of interaction. When the coherence length $L_{B}$ increases, the number of events with lower energy decreases due to the larger angular deflection. As expected, the curves corresponding to $\theta=40^{\circ}$ and $\theta=80^{\circ}$ have a higher maximum energy for larger distances.
@@
For $B=1\;\text{nG}$ (solid line), the peaks for $\theta=0^{\circ}$ result from the contribution of different primary particle families; just as in the previous scenario. Although protons dominate at lower energies, iron and intermediate nuclei prevail at higher energies. Secondary protons cannot reach Earth with energy above $\sim 6\times 10^{19}\;\text{eV}$ due to the large probability of photopion interaction. The tendency for different values of $L_{b}$ and $B$ is the same as that seen in the previous analysis of the primary studies of proton and nitrogen (figs. 2 and 4.
@@
The results for the magnetic field strength $B=10\;\text{nG}$ show an interesting feature in the modification factor for the misaligned jets. One may see that the curves corresponding to coherence lengths of $L_{B}=1\;\text{Mpc}$ and $L_{B}=10\;\text{Mpc}$ stand out. These higher values of $\xi$ are characterized by secondary protons that are sufficiently deflected to be able to reach Earth. This behavior is consistent with the nitrogen primary scenarios discussed previously for the same parameter values (fig. 4). Moreover, the magnetic horizon effect is also evident in this simulation. It is manifested as a suppression of the flux at lower energies when the source–Earth distance is $R=100\;\text{Mpc}$.
IV.3 The effect of the spectral index
@@
Our benchmark simulations were done for a spectral index $\gamma=1$. To investigate the dependence of the results on this parameter, we study another case: $\gamma=2$. We verified that the modification factor $\xi$ changes only slightly with the spectral index. As an illustration, Fig. 10 shows a comparison between $\xi$ as a function of energy, considering particles injected at sources with spectral indices $\gamma=1$ (red band) and $\gamma=2$ (yellow band). To improve the clarity of the plot, as the spectra overlap in some regions, the curves are plotted as bands, each with width $0.1$ along the vertical axis. All plots correspond to $L_{B}=0.1\;\text{Mpc}$ and $R=100\;\text{Mpc}$. Panel (a) depicts the case for $B=0.1\;\text{nG}$ and $\theta=0^{\circ}$, while panel (b) shows the results for $B=1\;\text{nG}$ and $\theta=40^{\circ}$. Simulations for $B=10\;\text{nG}$ and $\theta=80^{\circ}$ are presented in panel (c). It can be observed that, except for energies around $\sim 60$ EeV, corresponding to the energy threshold for photopion interactions, the behavior of the curves for both spectral indices is essentially the same, although a more pronounced Greisen-Zatsepin-Kuzmin (GZK) bump can be seen for $\gamma=1$, due to the larger contribution of more energetic particles compared to the case of $\gamma=2$.
@@
Figure 10: Modification factor obtained for primary protons, assuming an injected energy spectrum with spectral index $\gamma=1$ (red band) and $\gamma=2$ (yellow band). Each curve band has a $0.1$ width band along the vertical axis. All the results displayed correspond to a distance source-Earth ($R$) of $100$ Mpc, a $0.1$ Mpc coherence length ($L_{B}$) and a jet opening angle ($\Psi$) of $15^{\circ}$. The results are displayed for different combinations of magnetic field ($B$) and angle between the jet direction and the line of sight ($\theta$), from left to right: ($B$, $\theta$) $=$ ($0.1$ nG, $0^{\circ}$), ($1$ nG, $40^{\circ}$), ($10$ nG, $80^{\circ}$).
IV.4 The effect of the maximal rigidity
@@
Figure 11: Modification factors for nitrogen nuclei, obtained for a maximal rigidity of $R_{\text{max}}=10^{18.2}\;\text{V}$ (yellow band), and $R_{\text{max}}=10^{21}\;\text{V}$ (red), for different values of magnetic field ($B$) and various angles between the jet direction and the line of sight ($\theta$). The source is assumed to be distant 100 Mpc from Earth.
@@
In our benchmark scenario, we neglected the spectral cut-off of the sources. To study this effect separately, we present in Fig. 11 the modification factors obtained for nitrogen primaries, and for two different maximal rigidities, $R_{\text{max}}=10^{18.2}\;\text{eV}$ (red band) and $R_{\text{max}}=10^{21}\;\text{eV}$ (yellow band), for a fixed spectral index $\gamma=1$.
@@
There are noticeable differences in the modification factors, for reasons similar to what was discussed in IV.3: an interplay between emission geometry, magnetic-field strength, and interactions. The difference is, as expected, more pronounced at the higher energies. For stronger magnetic fields and larger viewing angles (right-most panel), the difference between the two $R_{\text{max}}$ case studied can be attributed to a change in composition — primary nitrogen nuclei dominate most of the flux below $E\lesssim 1\;\text{EeV}$, whereas the contribution of lighter nuclei is only dominant at higher energies.
IV.5 The emission geometry

### v2

@@ IV.1 Pure-proton scenario
where $\tau$ is the particle’s propagation time. Panel a corresponds to $\theta=0^{\circ}$, while panel b corresponds to $\theta=80^{\circ}$, both of which show the results for $B=10\;\text{nG}$ and $L_{B}=0.1\;\text{Mpc}$. The analysis of panel (a) reveals that most events reaching Earth have a very small elongation length, close to 1 Mpc. This indicates that the actual distance traveled by these particles is nearly equal to the source-Earth distance, resulting from minimal diffusion. As the arrival energy of these events decreases, the elongation length naturally increases. However, the counts for distances of $d\sim 10^{3}$ Mpc, which is of the same order of magnitude as the distance light would travel during the age of the Universe, are negligible. We conclude that the suppression observed in the lower energy part of the spectrum is not due to the high diffusion of these particles, which would cause the simulation time to reach the age of the Universe, but rather due to magnetic deflection. Unlike panel a, panel b, corresponding to $\theta=80^{\circ}$, indicates that a significantly larger number of protons propagated over distances greater than 1 Gpc (not shown in the plot, as the histogram represents the energy of detected particles). Consequently, their propagation times surpass the age of the Universe, preventing them from being detected on Earth.
Figure 3: Event count distribution as a function of the energy ($E$) and the trajectory elongation ($d$). Figure (a) presents the case where $\theta=0^{\circ}$, $B=10.0\;\text{nG}$, $L_{B}=0.1\;\text{Mpc}$ and $R=100\;\text{Mpc}$, while figure (b) shows the results for $\theta=80^{\circ}$, $B=10.0\;\text{nG}$, $L_{B}=0.1\;\text{Mpc}$ and $R=100\;\text{Mpc}$.
IV.2 Case of nuclei
@@
Figure 4: Modification factor obtained for primary nitrogen nuclei. The columns represent different source-Earth distances ($R$, and the rows vary according to the angle between the jet direction and the line of sight ($\theta$). The line styles indicate different magnetic field strengths ($B$) while the line colors correspond to different coherence lengths ($L_{B}$). Vertical lines mark the critical energy, with the color and line style matching the respective coherence length and magnetic field.
Figure 5: Energy distribution for primary nitrogen nuclei, for different arriving compositions. The figures show the results for the case where the jet direction is $\theta=0^{\circ}$, the magnetic strength considered is $B=0.1$ nG and $L_{B}=0.1$ Mpc. The columns vary in source-Earth distance, $R=10,\;40,\;100\;\text{Mpc}$ from left to right.
@@
Figure 4 shows the modification factor obtained after simulating $10^{5}$ nitrogen nuclei. The panels are organized in the same way as it was for protons, with rows corresponding to angles between the jet direction and the line of sight source-Earth and columns to different source-Earth distances. Due to the complexity introduced by nuclear fragmentation, we show energy distributions at Earth separated by particle type, for selected cases. We first analyze the results corresponding to $\theta=0^{\circ}$ (first row). The peak at the highest energies is significantly broader than in the case of proton simulations. This broadening arises because nitrogen nuclei fragment into multiple lighter elements with varying energies, producing a wider spread in the energy distribution of secondary protons arriving at Earth. Figure 5 shows the energy spectra for different nuclei, corresponding to the magenta dot-dashed lines in the first row of Fig. 4. These simulations were performed with an angle between the jet direction and the line of sight of $\theta=0^{\circ}$, magnetic field strength $B=0.1\;\text{nG}$, and coherence length $L_{B}=0.1\;\text{Mpc}$. From this figure, we observe that for a source distance of $R=10$ Mpc (first column), the energy range above $10^{20}$ eV is dominated by helium and protons resulting from nitrogen fragmentation. These secondary particles do not travel far enough to undergo photopion production or further nuclear fragmentation. This feature is absent in the spectra for $R=40$ Mpc (second column) and $R=100$ Mpc (third column), where the increased propagation distance enhances the probability of interactions with background photons. Additionally, the magenta dot-dashed curves in the first row of Fig. 4 show higher values of $\xi$ at $E\lesssim 6\times 10^{18}$ eV, due to the combination of nitrogen injected at the source with energies below the fragmentation threshold and an insufficient magnetic field strength and/or coherence length to cause larger deflections.
Figure 7: Energy distribution for primary nitrogen nuclei, for different arriving compositions. The panels show the results for the case where the magnetic strength considered is $B=10.0$ nG, the coherence length $L_{B}=10.0\;\text{Mpc}$, and $R=10$ Mpc. The left panel corresponds to a jet inclination of $\theta=40^{\circ}$ and the right one corresponds to $\theta=80^{\circ}$.
When considering other values of $L_{B}$, the main difference in the energy distribution per nuclear species is that, for higher values of $L_{B}$, the propagation regime becomes more diffusive at a given energy. As a result, the magnetic deflection enlarges and primary nuclei that would not disintegrate in the $L_{B}=0.1\;\text{Mpc}$ case may now undergo fragmentation as their path length increases. Despite this, the overall conclusions drawn from the analysis remain consistent with those obtained for $L_{B}=0.1\;\text{Mpc}$. As an example case, Fig. 6 shows a comparison between the energy spectra as a function of the coherence length $L_{B}$, where $L_{B}=0.1,\;1.0,\;10.0\;\text{Mpc}$ from left to right, $B$ is fixed on $1.0\;\text{nG}$ and $R$ on $40\;\text{Mpc}$.
@@
With respect to the cases of $\theta=40^{\circ}$ (second row), $\theta=80^{\circ}$ (third row) and the same magnetic field $B=0.1\;\text{nG}$ (dot-dashed line) discussed above, the curves are dominated by nitrogen emitted with energies below the fragmentation energy threshold. It may be reasonable to assume that the secondary particles produced from nitrogen photodisintegration do not experience enough deflection to reach Earth from misaligned jets with $\theta=40^{\circ}$ or $\theta=80^{\circ}$ in the presence of this magnetic field strength. One can also observe that as the source-Earth distance increases, more energetic particles have higher probability of arriving at Earth for the same parameter space due to the large angular deflection. As a consequence, the maximum energy observed in the spectra is shifted to the right, as expected and discussed in the pure-proton scenario. For a similar reason, the increase of $R$ must reveal an increasing contribution of lighter lower-energy nuclei, although it is expected that the primary nuclei remain dominant in the spectra. Similar conclusions are obtained considering $\theta=80^{\circ}$.
Figure 9: Energy distribution separated by the particle composition for the simulation of iron nuclei as primaries. The figures show the results for the case where the jet direction is $\theta=0^{\circ}$, the magnetic strength considered is $B=0.1$ nG and $L_{B}=1.0$ Mpc. The columns vary in source-Earth distance ($R$), $R=10,\;40,\;100\;\text{Mpc}$ from left to right.
For $B=1\;\text{nG}$ (continuous line), the peak at the highest energies, observed at $\theta=0^{\circ}$ (first row), is narrower for $L_{B}=1\;\text{Mpc}$ and $L_{B}=10\;\text{Mpc}$ compared to the peak for $L_{B}=0.1\;\text{Mpc}$ as a consequence of the loss of lower-energy events, resulting from angular deflections greater than $\Psi/2$. For $\theta=40^{\circ}$ and $\theta=80^{\circ}$, when we compare the results with those corresponding to $B=0.1\text{nG}$ (dot-dashed lines), we observe that the maximum energy achieved on Earth also increases. This result is in agreement with the conclusions addressed earlier in the pure-proton scenario.
@@
Figure 8 shows the modification factor resulting after simulating $10^{5}$ iron nuclei as primary particles. We first analyze the results corresponding to $\theta=0^{\circ}$ and $B=0.1\;\text{nG}$. The energy spectra for each particle type for these parameters, assuming $L_{B}=1.0\;\text{Mpc}$, are displayed in Fig. 9. For $R=10\;\text{Mpc}$, the peak at $E\sim 10^{19}\;\text{eV}$ observed in Fig. 8 is dominated by protons as can be seen in Fig 9. In contrast, the peaks at $E\sim 10^{19.5}\;\text{eV}$ and $E\gtrsim 10^{20}\;\text{eV}$ are respectively dominated by helium and iron primaries. For $R=40\;\text{Mpc}$ and $R=100\;\text{Mpc}$, the helium peak approaches the proton peak, as a result of the higher probability of interaction. When the coherence length $L_{B}$ increases, the number of events with lower energy decreases due to the larger angular deflection. As expected, the curves corresponding to $\theta=40^{\circ}$ and $\theta=80^{\circ}$ have a higher maximum energy for larger distances.
@@
For $B=1\;\text{nG}$ (solid line), the peaks for $\theta=0^{\circ}$ result from the contribution of different primary particle families; just as in the previous scenario. Although protons dominate at lower energies, iron and intermediate nuclei prevail at higher energies. Secondary protons cannot reach Earth with energy above $\sim 6\times 10^{19}\;\text{eV}$ due to the large probability of photopion interaction. The tendency for different values of $L_{b}$ and $B$ is the same as that seen in the previous analysis of the primary studies of proton and nitrogen (Figs. 2 and 4).
@@
The results for the magnetic field strength $B=10\;\text{nG}$ show an interesting feature in the modification factor for the misaligned jets. One may see that the curves corresponding to coherence lengths of $L_{B}=1\;\text{Mpc}$ and $L_{B}=10\;\text{Mpc}$ stand out. These higher values of $\xi$ are characterized by secondary protons that are sufficiently deflected to be able to reach Earth. This behavior is consistent with the nitrogen primary scenarios discussed previously for the same parameter values (Fig. 4). Moreover, the magnetic horizon effect is also evident in this simulation. It is manifested as a suppression of the flux at lower energies when the source-Earth distance is $R=100\;\text{Mpc}$.
IV.3 Effect of the spectral index
@@
Our benchmark simulations were done for a spectral index $\gamma=1$. To investigate the dependence of the results on this parameter, we study another case: $\gamma=2$. We verified that the modification factor $\xi$ changes only slightly with the spectral index. As an illustration, Fig. 10 shows a comparison between $\xi$ as a function of energy, considering particles injected at sources with spectral indices $\gamma=1$ (magenta band) and $\gamma=2$ (green band). To improve the clarity of the plot, as the spectra overlap in some regions, the curves are plotted as bands, each with width $0.1$ along the vertical axis. All plots correspond to $L_{B}=0.1\;\text{Mpc}$ and $R=100\;\text{Mpc}$. Panel a depicts the case for $B=0.1\;\text{nG}$ and $\theta=0^{\circ}$, while panel b shows the results for $B=1\;\text{nG}$ and $\theta=40^{\circ}$. Simulations for $B=10\;\text{nG}$ and $\theta=80^{\circ}$ are presented in panel c. It can be observed that, except for energies around $\sim 60$ EeV, corresponding to the energy threshold for photopion interactions, the behavior of the curves for both spectral indices is essentially the same, although a more pronounced Greisen-Zatsepin-Kuzmin (GZK) bump can be seen for $\gamma=1$, due to the larger contribution of more energetic particles compared to the case of $\gamma=2$.
@@
Figure 10: Modification factor obtained for primary protons, assuming an injected energy spectrum with spectral index $\gamma=1$ (magenta band) and $\gamma=2$ (green band). Each curve band has a $0.1$ width band along the vertical axis. All the results displayed correspond to a distance source-Earth ($R$) of $100$ Mpc, a $0.1$ Mpc coherence length ($L_{B}$) and a jet opening angle ($\Psi$) of $15^{\circ}$. The results are displayed for different combinations of magnetic field ($B$) and angle between the jet direction and the line of sight ($\theta$), from left to right: ($B$, $\theta$) == ($0.1$ nG, $0^{\circ}$), (11 nG, $40^{\circ}$), ($10$ nG, $80^{\circ}$).
IV.4 Effect of the maximal rigidity
@@
Figure 11: Modification factors for nitrogen nuclei, obtained for a maximal rigidity of $R_{\text{max}}=10^{18.2}\;\text{V}$ (magenta band), and $R_{\text{max}}=10^{21}\;\text{V}$ (green band), for different values of magnetic field ($B$) and various angles between the jet direction and the line of sight ($\theta$). The source is assumed to be distant 100 Mpc from Earth.
@@
In our benchmark scenario, we neglected the spectral cutoff of the sources. To study this effect separately, we present in Fig. 11 the modification factors obtained for nitrogen primaries, and for two different maximal magnetic rigidities, $R_{\text{max}}=10^{18.2}\;{\color[rgb]{0,0,0}\text{V}}$ (magenta band) and $R_{\text{max}}=10^{21}\;{\color[rgb]{0,0,0}\text{V}}$ (green band), for a fixed spectral index $\gamma=1$.
The spectral cutoff, according definition in Eq. 5, is given by
$\color[rgb]{0,0,0}f_{\text{cut}}(E,Z)=\begin{cases}1,&E\leq E_{\text{max}}\\ \exp\left(1-\frac{E}{E_{\text{max}}}\right),&E>E_{\text{max}},\end{cases}$} (9)
with $E_{\text{max}}=R_{\text{max}}\;Ze$.
@@
There are noticeable differences in the modification factors, for reasons similar to what was discussed in IV.3: an interplay between emission geometry, magnetic-field strength, and interactions. The difference is, as expected, more pronounced at the higher energies. For stronger magnetic fields and larger viewing angles (rightmost panel), the difference between the two $R_{\text{max}}$ cases studied can be attributed to a change in composition — primary nitrogen nuclei dominate most of the flux below $E\lesssim 1\;\text{EeV}$, whereas the contribution of lighter nuclei is only dominant at higher energies.
IV.5 Emission geometry